OpenStudy (anonymous):

Find how many four digit numbers can be formed using the digits 5,6,7,8 and 9 if; (a) any of the digits can occur more than once

OpenStudy (anonymous):

\[4^5\]

OpenStudy (anonymous):

the answer is 625

OpenStudy (anonymous):

??

OpenStudy (anonymous):

oh sorry i wrote \[4^5\] it is \[5^4\]

OpenStudy (anonymous):

why is it to the power of 4

OpenStudy (anonymous):

you have 4 slots to fill ___ ____ ____ _____ 5 choices for first slot, 5 for second, 5 for third, 5 for fouth

OpenStudy (anonymous):

"counting principle" tells you the number of ways you can do it is \[5\times 5\times 5\times 5\]

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

yw i bet there is a b and c and d too

OpenStudy (anonymous):

I've figured the others out lol